# Causal Discovery in Financial Markets: A Framework for Nonstationary Time-Series Data

### v2

@@ 6 Conclusion and Future Work
In this paper, we developed a novel algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), which is nonparametric (meaning it can capture both linear and non-linear causal relations), can handle nonstationarity, and is able to detect both lagged and contemporaneous causalities. We showed using simulated data that, for effective use of CD-NOTS, one should use ParCorr for low data regimes (less than 200 data points) and KCIT or RCoT for higher data regimes.
We applied CD-NOTS to Fama-French factors and Apple’s returns, where we found that for small time ranges (approximately seven years), Apple’s returns are stationary conditioning on the factors. This justifies the usage of Fama-French factors as a method for returns attribution. However, over longer time ranges (approximately two decades), we find Apple’s returns are nonstationary even after accounting for the Fama-French factors, implying that there is a need to continually update the factor exposures of Apple through time.
@@
Besides our contribution in developing an algorithm that can handle nonstationarity, non-linear relationships, lagged and contemporaneous causalities, we showed the importance of these assumptions in finance. Specifically, many of the assumptions made in prior work (notably linearity, no lagged relationships, and stationarity) often do not hold in finance and thus should be used with caution. Hence, through weakening the assumptions required for our causal discovery algorithm to be used, we can trust more the causal relationships found by CD-NOTS and continue to explore the causal relationships in other financial datasets.
@@ Appendix A Conditional Independence Tests Hyperparameters
For all conditional independence tests, we use a threshold of 0.05 for the p-value.
@@ Appendix B Fama-French Factors Case Study with RCoT
We repeat the plots from our first case study (Section 5.1) using RCoT as opposed to KCIT. One empirical finding was that, due to the randomness of RCoT, many times there could be an inconsistency between the discovered graph and the CI tests, i.e., the graph shows conditional independences that are not found by the CI tests. We leave it to future work to find how to account for these forms on inconsistencies.
Refer to caption Refer to caption Refer to caption Refer to caption (a) Jan 2000 - Dec 2007 Refer to caption (b) Jan 2008 - Dec 2015 Refer to caption (c) Jan 2016 - Dec 2022
Figure 13: Results of CD-NOTS RCoT run on three non-overlapping time periods.
Refer to caption Refer to caption
@@ Appendix C Economic Data Case Study with KCIT and RCoT
Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
Figure 15: Results of CD-NOTS KCIT run on six countries’ economic variables.
Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
Figure 16: Results of CD-NOTS RCoT run on six countries’ economic variables.
Figure 17: Results of CD-NOTS RCoT run on the economic variables of multiple countries.
@@ D.1 Different CI Tests
Figure 18: F-score evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
Figure 19: Precision evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
Figure 20: Recall evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
Figure 21: Structural Hamming Distance evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
D.2 Comparison against PCMCI
Figure 23: Precision evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
Figure 24: Recall evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
Figure 25: Structural Hamming Distance evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.

### v3

@@ 5 Conclusion and Future Work
In this paper, we developed a novel algorithm, Constraint-based Causal Discovery from Nonstationary Time Series (CD-NOTS), which is nonparametric, can handle nonstationarity, and is able to detect both lagged and contemporaneous causalities. We showed using simulated data that, for effective use of CD-NOTS, one should use ParCorr for low data regimes (less than 200 data points) and KCIT or RCoT for higher data regimes.
We applied CD-NOTS to approximately two decades of Fama-French factors and Apple’s returns and found that Apple’s returns are nonstationary even after accounting for the Fama-French factors, implying that there is a need to continually update the factor exposures of Apple through time.
@@
Through our case studies, we show the importance of testing the assumptions of causal discovery algorithms. Specifically, we find in all of our case studies nonstationarity, non-linear relationships, and lagged and contemporaneous causal effects. Prior work often assumes some of these properties are not present in the data and thus would lead to spurious relationships and incorrect conclusions. Hence, our contribution improves over prior work by handling all of these scenarios.
@@ Appendix A Conditional Independence Tests Hyperparameters
For all conditional independence tests in our case studies, we use a threshold of 0.05 for the p-value.
@@ Appendix B Fama-French Factors Case Study with RCoT
We repeat the plots from our first case study (Section 4.1) using RCoT as opposed to KCIT. One empirical finding is that, due to the randomness of RCoT, many times there could be an inconsistency between the discovered graph and the CI tests, i.e., the graph shows conditional independences that are not found by the CI tests. We leave it to future work to find how to account for these forms of inconsistencies.
@@ Appendix C Economic Data Case Study with KCIT and RCoT
Our synthetic data experiments indicate the effectiveness of ParCorr when data is limited, as observed in our country-level experiments. However, in this section, we present results obtained using KCIT and RCoT. Notably, these results appear to lack robustness to specific test used, emphasizing the importance of analyzing countries collectively. Upon comparing KCIT with RCoT, we note several similarities in edges, barring a few exceptions. For instance, RCoT suggests that Unemployment and PPI are stationary, whereas CPI lacks a one-lag relationship with itself. Additionally, it indicates an additional edge from Unemployment in the previous month to the current month’s CPI.
(a) U.S.
(b) Canada
(c) Japan
(d) France
(e) United Kingdom
(f) Italy
Figure 17: Results of CD-NOTS KCIT ran on six countries’ economic variables.
(a) U.S.
(b) Canada
(c) Japan
(d) France
(e) United Kingdom
(f) Italy
Figure 18: Results of CD-NOTS RCoT ran on six countries’ economic variables.
@@ Appendix D Detailed Simulation Results
Further findings on simulated data includes an evaluation of precision, recall555precision$=\frac{tp}{tp+fp}\,$, recall$=\frac{tp}{tp+fn}$ where $tp=$true positive, $fp=$false positive and $fn=$ false negative. and Structural Hamming Distance (SHD). Note that while precision, recall, and F-score aim for maximization, the objective for SHD is minimization.
Figure 19: Results of CD-NOTS RCoT ran on the economic variables of multiple countries.
@@ D.1 Different CI Tests
At 300 data points, KCIT-SW demonstrates the highest recall but substantially lower precision, resulting in the poorest F-score within the plot. Conversely, in the context of 1000 data points, recall remains consistent across all methods, with discrepancies in F-score attributed to variations in precision. Notably, with 50 data points, the precision curve exhibits an unexpected upward trend, contrary to the anticipated decrease in precision (and recall) as the node count increases. Interestingly, despite these nuances, SHD yields a performance ranking similar to that of F-score across the evaluated scenarios.
Figure 20: Precision evaluation for CD-NOTS with different CI tests, tested out on different simulated datasets.
Figure 21: Recall evaluation for CD-NOTS with different CI tests, tested out on different simulated datasets.
Figure 22: SHD evaluation for CD-NOTS with different CI tests, tested out on different simulated datasets.
D.2 Effect of $\alpha$
Note that the significance level $\alpha$ controls the false positive rate. i.e., when employing a conditional independence test, if the data is truly independent, the test will fail $\alpha$ percent of the time, yet it does not explicitly address false negatives. This phenomenon becomes evident when observing precision plots, where the optimal precision corresponds to the lowest $\alpha$, indicating the lowest false positive rate. Conversely, recall operates inversely, as a decrease in false positives generally coincides with an increase in false negatives. It is this balance that the F-score metric achieves, although it is worth noting that F-score is not the sole method to combine the two metrics together.
Regarding SHD, it presents another perspective on balancing false positives and false negatives. Unlike F-score, SHD typically advocates for an alpha of 0.01, except for specific scenarios, such as when dealing with 50 and 300 data points in RCoT and CMIknn for 300 data points with larger graphs.
(a) ParCorr
(b) CMIKnn
(c) KCIT
(d) RCoT
Figure 23: Effect of $\alpha$ on Precision
(a) ParCorr
(b) CMIKnn
(c) KCIT
(d) RCoT
Figure 24: Effect of $\alpha$ on Recall
(a) ParCorr
(b) CMIKnn
(c) KCIT
(d) RCoT
Figure 25: Effect of $\alpha$ on SHD
D.3 Comparison against PCMCI
CD-NOTS consistently outperforms PCMCI in terms of SHD, similar to its dominance in the F-score metric. While PCMCI exhibits better recall than CD-NOTS, it does so at the cost of significantly lower precision. This trade-off explains why CD-NOTS achieves better performance in both F-score and SHD metrics.
Figure 26: Precision evaluation, comparing CD-NOTS and PCMCI, tested out on different simulated datasets.
Figure 27: Recall evaluation, comparing CD-NOTS and PCMCI, tested out on different simulated datasets.
Figure 28: SHD evaluation, comparing CD-NOTS and PCMCI, tested out on different simulated datasets.